OpenStudy (anonymous):

3. Two ocean beaches are being affected by erosion. The table shows the width, in feet, of each beach at high tide measured where 1995 is represented by year 0.

OpenStudy (anonymous):

A. Describe the patterns shown by the erosion data measurements shown for each of the beaches in the table. B. Between which years will the beaches have approximately the same width? C. Assuming these rates remain constant, what can you do to get a better approximation of when the two beaches will have the same width?

OpenStudy (anonymous):

2.5x – 10.5 = 64(0.5x) 2.5x - 10.5 = 32x 2.5x - 32x = 10.5 -29.5x = 10.5 x = -10.5/29.5

OpenStudy (anonymous):

we are not aloud to give you the answer on here but I don't know how to walk you through it but I do know the answer!

OpenStudy (anonymous):

@alyssa23 yes please walk me through

OpenStudy (anonymous):

Give me one sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im taking a test so I will be right back

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im need to finish my test so hold on

OpenStudy (anonymous):

okay thankyou

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

1.Complete the table below to solve the equation 2.5x – 10.5 = 64(0.5x).? x f(x) = 2.5x – 10.5 g(x) = 64(0.5x) 2 -5.5 64 3 -2.5 96 4 -0.5 128 I think there is no solution for x > 0 0 -10.5 0 0.1 0.2 -0.3556 -11.3898 -11.3898 Let us try algebra. 2.5x – 10.5 = 64(0.5x) = 32x or 29.5x = -10.5 or x = -10.5/29.5 = -.03556

OpenStudy (anonymous):

Western Beach is losing sand, Dunes beach is gaining sand. Between 11 and 12 years Plot the data against time and draw a curve that goes through the data. Where the two lines cross is the time when both beaches are the same width.